Calculate the pH of 0.20 M benzoic acid and 0.10 M sodium benzoate solution.
Ka (HC7H5O2) = 6.5 × 10-5.
1
Expert's answer
2016-03-18T15:48:05-0400
The formula of the pH of buffer solution:
pH = pKa(acid)+lg(C(salt)/C(acid))
pKa(HC7H5O2) = -lg(6.5 × 10-5) = 4.187
pH = 4.187 +lg((0.10 M)/(0.20 M)) = 4.187 – 0.301 = 3.886

Answer:
The pH of this solution is 3.886.